What is the missing coefficient in the equation given that p(2) = 128?
p(x) = _________ x5

Answer by stanbon(75887)

What is the missing coefficient in the equation given that p(2) = 128?
p(x) = x^5
---------
128 = x^5
----------------
x = 128^(1/5)
x = 2
-------------
Cheers,
Stan H.
-------------